OpenStudy (anonymous):

(xy + 3) (xy - 1) would that be x^2y^2 + 2xy - 3?

jagr2713 (jagr2713):

correct

OpenStudy (anonymous):

thank you

jagr2713 (jagr2713):

Anytime

OpenStudy (anonymous):

how would i do this..? (x - 3) (x^2 - 2x + 4) ?

jagr2713 (jagr2713):

any ideas ?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

well i was thinking i could do x times 2x and get 2x^2 and then 3 times 4...? im not sure

OpenStudy (anonymous):

@sleepyjess can u help?

OpenStudy (sleepyjess):

This is just a longer version of FOIL. Give me just a minute :)

OpenStudy (anonymous):

thank you

jagr2713 (jagr2713):

let me let jess do this she needs 99!!!!!!

OpenStudy (sleepyjess):

(x - 3)(x^2 - 2x + 4) x*x^2 -3*x^2 now we are done with x^2 x*-2x -3*-2x now we are done with -2x x*4 -3*4 this is all :)

OpenStudy (anonymous):

im lost..

OpenStudy (anonymous):

out of that i got -3x^2 - 6x^2...? idk

OpenStudy (sleepyjess):

x*x^2 = x^3 -3*x^2 = -3x^2 x*-2x = -2x^2 -3*-2x = 6x x*4 = 4x -3*4 = -12 Now put them all together into one equation and combine like terms

OpenStudy (anonymous):

x^3 - 5x^2 + 10x - 12 0r....-x^3 - 3x^2 + x - 12?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

the first one is correct

OpenStudy (anonymous):

thx

OpenStudy (sleepyjess):

no problemo :)

OpenStudy (anonymous):

would (x+4)(-x^2+x-3) be -x^3-3x^2+x-12?

OpenStudy (sleepyjess):

x*-x^2 = -x^3 4*-x^2 = -4x^2 x*x = x^2 4*x = 4x x(-3 = -3x 4*-3 = -12 -x^3 - 3x^2 + x -12

OpenStudy (sleepyjess):

yes

OpenStudy (anonymous):

okay just one more

OpenStudy (sleepyjess):

okie ^_^

OpenStudy (anonymous):

(2x-1)(-x^2-x+4) would that be 2x^3-10x^2-26x-14 or -2x^3-2x^2+9x-5?

OpenStudy (sleepyjess):

2x*-x^2 = -2x^3 -1*-x^2 = x^2 2x*-x = -2x^2 -1*-x = x 2x*4 = 8x -1*4 = -8 -2x^3 + x^2 - 2x^2 + x + 8x - 8 -2x^3 - x^2 + 9x - 8

OpenStudy (sleepyjess):

ah, -1*4 does not equal -8, should be -4

OpenStudy (anonymous):

yeah im sorry i have answer choices and i looked at them and i typed one of them wrong

OpenStudy (anonymous):

i mest up with the 4

OpenStudy (anonymous):

so it would be -2x^3 - 2x^2 + 9x - 4?

OpenStudy (sleepyjess):

yes

OpenStudy (anonymous):

Thank you so much, you helped me with everything and that means a lot, so thank you :)

OpenStudy (sleepyjess):

no problem! :)